Antonio received 40 ISOs (each option gives him the right to purchase 20 shares of Zorro stock for $3 per share) at the time he started working for Zorro Corporation six years ago. Zorro’s stock price was $3 per share at the time. Now that Zorro’s stock price is $50 per share, Antonio intends to exercise all of his options and immediately sell all the shares he receives from the options exercise. (Enter all amounts as positive values. Leave no answers blank. Enter zero if applicable.)

a. What are Antonio’s taxes due on the grant date, the exercise date, and the date the shares are sold, assuming his ordinary marginal rate is 32 percent and his long-term capital gains rate is 15 percent?
Answer #1

Shares Acquired = No. of ISOs * Shares in each ISO = 40 * 20 = 800 shares

Cash Needed to Exercise = Shares Acquired * Exercise Price = 800 * $3 = $2,400

Amount Realized = Shares Acquired * Market Price = 800 * $50 = $40,000

Ordinary Income = Amount Realized - Cash Needed to Exercise = $40,000 - $2,400 = $37,600

Tax Paid in year of sale = Ordinary Income * Ordinary Marginal Tax Rate

= $37,600 * 0.32 = $12,032